OpenStudy (anonymous):

1 × 3^1 + 2 × 3^2 + 3 × 3^3 + … + n × 3^n = 3/4 (35*3^18 + 1) What is the value of n?

OpenStudy (anonymous):

I cant think of a short way to do this problem. I have a pretty lengthy solution that shows n = 18

OpenStudy (anonymous):

This is the sum of the terms of a sequence of the form \[U(n)=aq^n\] With a=n and q=3. The sum n terms is given by \[s_{n}=a \frac{1-q^n}{1-q}\] I'd try starting with that no?

OpenStudy (anonymous):

The way I did it was: First use the geometric series formula: \[1+x+x^2+\cdots+x^n = \frac{x^{n+1}-1}{x-1}\] Take the derivative of both sides: \[1+2x+3x^2+\cdots+nx^{n-1} = \frac{nx^{n+1}-(n+1)x^n+1}{(x-1)^2}\]Multiply both sides by x: \[x+2x^2+3x^3+\cdots+nx^n = \frac{nx^{n+2}-(n+1)x^{n+1}+x}{(x-1)^2}\] Notice if you let x = 3, the left hand side becomes the same as the left hand side of the problem that was asked. So plug x = 3 into the right hand side, and simplify, then set equal to the right hand side of the problem. You will end up with n = 18. Like i said, i dont know if there is a shorter way.

OpenStudy (anonymous):

When x = 3 you get: \[\frac{n(3)^{n+2}-(n+1)3^{n+1}+3}{(3-1)^2} = \frac{9n(3)^n-3(n+1)(3)^n+3}{4} \] \[\frac{(9n-3n-3)(3)^n+3}{4} = \frac{(6n-3)3^n+3}{4} = \frac{3}{4}((2n-1)3^n+1) \] Setting that equal to the right hand side of the problem we obtain: \[\frac{3}{4}((2n-1)3^n+1) = \frac{3}{4}(35(3)^n+1)\] Thus we have: \[2n-1 = 35, n = 18\]

OpenStudy (dumbcow):

very nice..i never thought of taking the derivative. well the easy way of solving this is using a program or excel and compute it numerically since we know what the sum is. however, mathematically speaking i found it can be broken down into a series of smaller geometric series \[3^{1} + (2)3^{2}...+(n)3^{n} = (3^{1}+3^{2}...+3^{n}) + (3^{2}+3^{3}...+3^{n})...+3^{n}\] using formula for geometric sum for each one \[\frac{(3^{n+1}-3)+(3^{n+1}-3^{2})...+(3^{n+1}-3^{n})}{3-1}\] which can reduce to \[\frac{3[3^{n}(2n-1)]}{4} \]

OpenStudy (anonymous):

wow that is also very nice, im going to take not of that :)